Tohban Report 2017-03-15

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 8 March 2017
End Date: 15 March 2017
Tohban: Milo Buitrago-Casas
Tohban email:
Next Tohban: TBD
List all reports


Solar Activity

Solar activity was very low this week. Currently, there is a very small active region on disk but with very low chances to host a flare. Solar activity is expected to continue at very low levels with a chance for B-class flares in a coming couple of days based on recent region trends and lack of activity (continuity).

The RHESSI Tohban Flare counter reported no-flares this past week.

Memory Management

Memory was ranging between 30% and 40% at the end of last week. But with the resumption of the BGS passes on Friday 10 around noon (PST), the SSR filling has been uniformly dropping down.

Memory is steady at ~25% at end of pass daily set.

Decimation was normal/vigorous.

Spacecraft Status

D3 and D6 are on. Cryocooler power is running "nominal" at 75W. According to Brian's plots, the cold plate temperature has risen from ~130 K in June 2016 to ~150 K now or an increase of 20 K in 9 months or 2.2 K per month or 0.07 K per day in Albert’s units. The accelerometer reading showed large fluctuations when the power was at 85 watts from September to December 2016 but has been lower and more stable since the power was reduced to 75 watts although it has been slowly increasing since January with the power at 75 watts.

Data Gaps

GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-14T12:05:00.000 -- 2017-03-14T12:10:00.000       300.00000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-14T09:55:00.000 -- 2017-03-15T00:00:00.000       50700.000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-13T01:40:00.000 -- 2017-03-13T01:45:00.000       300.00000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-12T20:10:00.000 -- 2017-03-12T21:55:00.000       6300.0000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-12T20:10:00.000 -- 2017-03-12T21:50:00.000       6000.0000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-09T18:10:00.000 -- 2017-03-09T18:15:00.000       300.00000
2017-03-09T18:35:00.000 -- 2017-03-09T18:40:00.000       300.00000
2017-03-09T19:20:00.000 -- 2017-03-09T19:25:00.000       300.00000
2017-03-09T19:30:00.000 -- 2017-03-09T19:35:00.000       300.00000
2017-03-09T19:40:00.000 -- 2017-03-09T19:45:00.000       300.00000
2017-03-09T20:25:00.000 -- 2017-03-09T20:30:00.000       300.00000
GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-08T16:20:00.000 -- 2017-03-08T17:00:00.000       2400.0000

Detector issues

Detectors D3 and D6 are on and working. Some high counts at low energies that cause the high SSR filling rate. We should find a good compromise among increasing the front slow threshold, increasing the high voltage, and keep low energies data.

There were no maneuvers on the detectors during last week.


Tohban report from last week needs to be upload.

Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) plus/minus 4 minutes from daylight
Night time data (rears) plus/minus 4 minutes from daylight
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Commanded out
Detector problems? See notes above.
Personal tools